# Careless Whisper: Speech-to-Text Hallucination Harms

### v1

@@ 3. Results
Across both April and May 2023 transcriptions generated, we identify 312 instances of hallucinations (as some audio segments resulted in hallucinations only in April but not May, vice versa, and both times). We categorize these hallucinations in Figure 0(a), and find that among the 312 hallucinated transcriptions, 19% include harms perpetuating violence, 13% include harms of incorrect association, and 8% include harms of fraudulent video-based authority or phishing. Across all hallucinated transcriptions, 38% contain at least one of these types of harms. Hallucinated transcriptions expand our notion of AI harms in very concrete ways since they can cause real-world harms (Barocas et al., 2017): violent content can trigger victims of abuse, representational harm can occur due to incorrect association, and fraudulent authority or phishing attempts can cause financial loss.
We further tested Whisper as of December 2023 on the aforementioned set of 187 audio segments that reliably resulted in hallucinations in Spring 2023. While we found that many of the aforementioned examples of hallucinations were resolved, it remained the case that 12 out of 187 audio segments continued to result in hallucinations (9/92 aphasia audio files and 3/95 control audio files). This significant improvement is likely the result of a late-November 2023 update to Whisper that explicitly skips periods of silence in audio to reduce hallucinations. [^5] https://github.com/openai/whisper/pull/1838 Next, we re-ran a random sample of 250 audio segments (among the original 13,140 segments) through Whisper in December 2023 to compare to the Spring 2023 runs. [^6] This sample included 38 control segments and 212 aphasia segments. Among this random sample, we found only 1 out of 250 instances where hallucinations were introduced in the December 2023 run but not in either Spring 2023 run. This example showcases the Made-up Name harm for an aphasia speaker, where the ground truth was: “[RedactedName] and I, [RedactedName] did”, and the December 2023 Whisper transcription was: “Um, [RedactedName] and I, um, [RedactedName], um, did, um… Actually, I wanted to ask why would my friends pair them up to go see Dorothy Maxwell. Um, um… Um… Um…’. On our dataset, these trends suggest that Whisper is improving at hallucination reduction—but still regularly and reproducibly hallucinates.
(a) While some hallucinated text could be considered innocuous despite being incorrect, a concerning 38% of the hallucinated text falls under one of three identified harmful categories.
@@
Figure 1. Hallucinations are more common for speakers with aphasia than without, and can cause harm by nature of perpetuating violence, incorrect identifications or associations, and false authority or phishing.
@@ 3.1. Limitations
Our work compares a relatively small set of aphasia speakers to control group speakers in a setting where they are being asked a standard slate of interview questions. If a broader set of topics were to be discussed, it is possible that the scope of hallucinations would be widened. Furthermore, we note that our tabulation of hallucinations is likely an undercount due to the initial restriction on yielding different outcomes between the April and May 2023 API runs. In the case where both April and May 2023 transcriptions included hallucinations, these would not be counted in our above tabulation.
Finally, there is one category of hallucinations that we did not tabulate in the above counts due to the additional resources necessary to determine whether they would indeed qualify as a hallucination: the appearance of other languages. For example, Whisper is prone to generating non-English transcriptions even when provided an argument indicating that the target language is English. In some cases, this is a bonus: for individuals who code switch, transcriptions can be provided accurately in the correct language. However, in other cases, the non-target language is simply being hallucinated. See Figure 2 for an example of each case: while the example in Ukranian is a faithful transcription of the true speech, the example in Chinese shows telltale signs of hallucination (e.g., repeated copies of a phrase), and ends with four characters translating to “thanks for watching”—an example of the harm of false video-based authority or phishing.
Figure 2. Two examples of control speakers whose Whisper transcriptions from December 2023 included non-English text, despite the API setting for language being set to English. The first example is not a hallucination, whereas the second example is hallucinated.
@@ 4.1. Generative AI in speech-to-text modeling
Our first hypothesis involves the fact that most advanced speech-to-text technology combines language modeling and audio modeling. The basis for OpenAI’s language model could be similar to hallucination-prone ChatGPT; in fact, Whisper even allows user prompting “for correcting specific words or acronyms that the model often misrecognizes in the audio” (OpenAI, 2023b). In our experiment, we did not input any prompts, and set the sampling temperature parameter (which controls the randomness of responses) to the default of 0, which should yield minimally random transcriptions.
Despite this, our experiment yielded highly non-deterministic hallucinations – perhaps implying that Whisper’s over-reliance on OpenAI’s modern language modeling is what leads to hallucinations. Notably, we found no evidence of hallucinations in competing speech recognition systems such as Google Speech-to-Text (tested in April 2023) or the latest Google Chirp model (tested in December 2023): we identified exactly 0 comparable hallucination concerns (as defined above) from Google’s products out of the 187 identified audio segments. We similarly identified exactly 0 comparable hallucination concerns among the same 187 audio segments tested on Amazon, Microsoft, AssemblyAI, and RevAI speech-to-text services tested in January 2024. This could indicate that generative language models such as PaLM2 (underlying Google Bard) are not being used in a similar manner in this speech-to-text systems.
@@ 4.2. Speech patterns likely to yield hallucinations
In our experiment, we compare audio from American English speakers with aphasia (5,335 audio segments) to audio spoken by the control group (i.e., those speaking “standard” American English without speech disorders, in 7,805 audio segments). First, we find that Whisper transcription of aphasia speakers yields significantly more hallucinations (p=0.019) than for the control group. Specifically, 1.7% aphasia audio segments yielded hallucinations, versus only 1.2% of control group audio segments (per Figure 0(b)).
@@
Next, we seek to determine if longer pauses (disfluencies) in aphasia speech could contribute to the higher rate of hallucinations. We quantify these disfluencies by using VAD (Voice Activity Detection) to measure the pauses or weak signals in audio files. First, we calculate the “non-vocal duration” of each audio file in milliseconds using PyAnnote [^7] We also run VAD analysis via PyTorch Hub using Silero (Team, 2021) and find comparable results; see the Appendix for more details. (Bredin and Laurent, 2021; Plaquet and Bredin, 2023). Then, we can calculate the share of the audio segment that is non-vocal—i.e., the percentage of audio that does not consist of any (or only very weakly detected) human speech. Per Figure 3 there are larger non-vocal shares of total durations for aphasia speakers relative to control speakers (41% versus 15%, respectively; p-value ¡ 2.2e-16). Furthermore, we can compare the non-vocal durations within each group between audio files that yielded hallucinations and audio files that did not yield hallucinations. Audio files yielding hallucinations tended to have higher non-vocal shares of durations (29% versus 26% of total durations). [^8] In Appendix Figure 4(b), we also see that (as expected), there are significantly longer non-vocal durations (in seconds) for aphasia speakers relative to control speakers (6.8 seconds versus 1.3 seconds, respectively; p-value ¡ 2.2e-16). The non-vocal durations in audio files were also significantly higher in segments yielding hallucinations relative to those that did not (5.3 seconds versus 3.5 seconds).
@@
Figure 3. Speakers with aphasia had audio files with significantly longer shares of non-vocal sounds (i.e., PyAnnote non-vocal duration divided by total duration) relative to their control speaker counterparts. Furthermore, non-vocal shares of audio files were significantly higher for files with Whisper hallucinations as opposed to files that did not yield hallucinations. Mean non-vocal shares for aphasia speakers with hallucinations, aphasia speakers without hallucinations, control speakers with hallucinations, and control speakers without hallucinations are: 42.4%, 40.6%, 16.2%, and 15.4%, respectively.
Finally, we train a logistic regression on the audio segments with corresponding speaker demographic information to predict whether an output will be hallucinated based on speaker demographics and audio characteristics; per Table 2, we find that the most significant attributes positively corresponding to a Whisper hallucination are speaking with aphasia, having a longer non-vocal duration, and speaking with more words. This is consistent with our hypothesis on disfluencies, and is robust to other regression specifications, including on Mahalanobis-matched subsets of audio segments (see Appendix).
Table 2. By fitting a logistic regression conditioned on speaker demographics and audio segment attributes, we find that being a speaker with aphasia, having longer non-vocal periods in an audio recording, and speaking with more average words (in the ground truth transcription) per segment duration, all tend to result in a significantly higher likelihood of a hallucinated Whisper transcription.
@@
Has Aphasia ${}^{**}$ (0.257)
Non-vocal Duration (seconds) ${}^{***}$ (0.015)
Average Word Speed ${}^{***}$ (0.128)
Is Female --0.019 (0.168)
Age 0.045 (0.043)
@@
African American --0.174 (0.468)
Other Race --0.311 (0.518)
Years of Education --0.048 (0.033)
English is First Language --0.368 (1.037)
No Vision Loss 0.418 (1.023)
No Hearing Loss 0.442 (1.024)
Constant ${}^{***}$ (2.169)
@@
Log Likelihood --785.562
Akaike Inf. Crit. 1,597.125
Note: $*{}^{*}$p$<$0.1; ${}^{**}$p$<$0.05; ${}^{***}$p$<$0.01
@@ 5. Implications and Future Work
Our work demonstrates that there are serious concerns regarding Whisper’s inaccuracy due to unpredictable hallucinations. It also demonstrates that there are significant issues with the reproducibility of transcriptions in light of non-deterministic hallucinations. Most importantly, however, there are implications for understanding the harms and biases arising from the disparate impacts on subpopulations whose speech is likely to yield hallucinations (Papakyriakopoulos et al., 2023). As we demonstrate above, there are specific harms that are produced by hallucinations for individuals with aphasia: the perpetuation of violence, the incorrect identification or association with someone, and harms of false authority or phishing. All there categories can have serious implications for the aphasic speaker. Hallucinated violent speech may be flagged by institutions in power, potentially leading to a premature dismissal of an applicant in a hiring process, or an individual interacting with a government service chat bot. Incorrect identification or association, similarly, could have severe impacts on an aphasic speaker, for example in the context of legal proceedings or the interaction with insurance companies. Lastly, authority or phishing hallucinations could misclassify an aphasic speaker as spam or facilitate future cyberattacks.
These three types of harms can be interpreted as both allocative and representational harms (Barocas et al., 2023), i.e., harms that are caused when a system withholds certain groups an opportunity or a resource and harms that occur when systems reinforce the subordination of some groups along the lines of identity (including race, class, gender, etc.). This is because hallucinations can affect how individuals with aphasia can gain access to opportunities, such as jobs, and resources, such as information about government services, and because hallucinations build on existing (harmful) stereotypes and reinforce them.
@@
Future work in this area should, therefore, examine the intersectional nature of hallucination bias and harms. We suggest that researchers may use our approach for identifying and categorizing hallucination harms, i.e., explicitly account for the type of technology (in this case: generative AI), the particular application (in this case: speech-to-text transcription), and the particular subpopulation (in this case: individuals with aphasia). This approach can avoid generalizations of AI harms and allow for more specific AI harms classifications that may lead to targeted mitigation strategies.
The primary outcome of this paper focuses on the existence of hallucinations, including hallucinations of different types. However, future work can and should be done using the Word Error Rate (WER), a standard metric of accuracy for speech-to-text. Prior work has found that WER is consistently worse for populations with health concerns, such as individuals with dysphonia (Lopez et al., 2023). Definitionally, hallucinations will also cause the WER metric to be worse. Forthcoming work regards comparing WERs across different speech-to-text services, wherein we find that Whisper performs better on average relative to its peers. This leads to the question: should users accept use of an overall-accurate system, if it occasionally makes massive transcription errors in the form of hallucinations?

### v2

@@ 3. Results
Across both April and May 2023 transcriptions generated from the 187 audio segments yielding hallucinations, we identify 312 transcriptions containing hallucinations (as some audio segments resulted in hallucinations only in April but not May, vice versa, and both times). On average, we found that 1.4% of transcriptions in our dataset yielded hallucinations. We categorize these hallucinations in Figure 1(a), and find that among the 312 hallucinated transcriptions, 19% include harms perpetuating violence, 13% include harms of inaccurate associations, and 8% include harms of false authority. Across all hallucinated transcriptions, 38% contain at least one of these types of harms. Hallucinated transcriptions expand our notion of AI harms in very concrete ways since they can cause real-world harms (Barocas et al., 2017): violent content can trigger victims of abuse, incorrect association can cause representational harm, and fraudulent authority or phishing attempts can cause financial loss.
We further tested Whisper as of December 2023 on the aforementioned set of 187 audio segments that reliably resulted in hallucinations in April or May 2023. While we found that many of the aforementioned examples of hallucinations were resolved, it remained the case that 12 out of 187 audio segments continued to result in hallucinations (9/92 aphasia audio files and 3/95 control audio files). This significant improvement is likely the result of late-November 2023 updates to Whisper. [^5] One such improvement (that we do not use in our experiments) is the introduction of a parameter allowing the user to explicitly skip specific durations of silence in the beginning of their audio files, in order to reduce hallucinations. https://github.com/openai/whisper/pull/1838 Next, we re-ran a random sample of 250 audio segments (among the original 13,140 segments) through Whisper in December 2023 to compare to the April/May 2023 runs. [^6] This sample included 38 control segments and 212 aphasia segments. Among this random sample, we found only 1 out of 250 instances where hallucinations were introduced in the December 2023 run but not in either April or May 2023 run. This example showcases the Made-up Names harm for an aphasia speaker, where the ground truth was: “[RedactedName] and I, [RedactedName] did”, and the December 2023 Whisper transcription was: “Um, [RedactedName] and I, um, [RedactedName], um, did, um… Actually, I wanted to ask why would my friends pair them up to go see Dorothy Maxwell. Um, um… Um… Um…’. In our dataset, these trends suggest that Whisper is improving at hallucination reduction—but still regularly and reproducibly hallucinates.
(a) While some hallucinated text could be considered innocuous despite being
incorrect, a concerning 38% of the hallucinated text falls under one of three
identified harmful categories.
@@
Figure 1. Hallucinations are more common for speakers with aphasia than without, and can cause harm by nature of perpetuating violence, inaccurate associations, and false authority.
@@ 3.1. Limitations
Our work compares a relatively small set of aphasia speakers to control group speakers in a setting where they are being asked a standard slate of interview questions. If a broader set of topics were to be discussed, it is possible that the scope of hallucinations would be widened. Furthermore, we note that our tabulation of hallucinations is likely an undercount due to the initial restriction on non-deterministically yielding different outcomes between the April and May 2023 API runs. In cases where both April and May 2023 transcriptions included the same hallucinations, these hallucinations would not be counted in our above tabulation.
Finally, there is one category of hallucinations that we did not tabulate in the above counts due to the additional resources necessary to determine whether they would indeed qualify as a hallucination: the appearance of other languages. For example, Whisper is prone to generating non-English transcriptions even when provided an argument indicating that the target language is English. In some cases, this is a bonus: for individuals who code switch, transcriptions can be provided accurately in the correct language. However, in other cases, the non-target language is simply being hallucinated. See Figure 2 for an example of each case: while the example in Ukranian is a faithful transcription of the true speech, the example in Chinese shows telltale signs of hallucination (e.g., repeated copies of a phrase), and ends with four characters translating to “thanks for watching”—an example of the harm of false authority.
Figure 2. Two examples of control speakers whose Whisper transcriptions from December 2023 included non-English text, despite the API setting for language being set to English. The first example is not a hallucination, whereas the second example is hallucinated (involving a repeating loop, and displaying a harm of false authority: thanking.)
@@ 4.1. Generative AI in speech-to-text modeling
Our first hypothesis involves Whisper’s speech-to-text modeling choices. Most advanced speech-to-text systems now use a single end-to-end model rather than combining separate acoustic and language models. OpenAI’s Whisper model could involve similar technology to hallucination-prone language model ChatGPT; in fact, Whisper even allows user prompting “for correcting specific words or acronyms that the model often misrecognizes in the audio” (OpenAI, 2023b). In our experiment, we did not input any prompts, and set the sampling temperature parameter (which controls the randomness of responses) to the default of 0, which should yield minimally random transcriptions. Despite this, our experiment yielded highly non-deterministic hallucinations—perhaps implying that Whisper’s over-reliance on OpenAI’s language modeling advancements is what leads to hallucinations.
Furthermore, the datasets ingested by OpenAI systems can help to explain the existence of the “False Authority” class of harms—and in particular, “Youtuber” language in hallucinations. Contemporaneous reporting (Metz et al., 2024) revealed that Whisper was used to transcribe over a million hours of YouTube videos in order to train GPT-4. This is consistent with the high volumes of “False Authority” harms that we identified.
Notably, we found no evidence of hallucinations in competing speech recognition systems such as Google Speech-to-Text (tested in April 2023) or the latest Google Chirp model (tested in December 2023): we identified exactly 0 comparable hallucination concerns (as defined above) from Google’s products out of the 187 identified audio segments. We similarly identified exactly 0 comparable hallucination concerns among the same 187 audio segments from Amazon, Microsoft, AssemblyAI, and RevAI speech-to-text services (tested in January 2024). This could indicate that advancements in generative language models such as PaLM2 (underlying Google Bard) were not being used in a similar manner in competing speech-to-text systems. As such, we believe hallucinations to currently be an OpenAI-specific concern.
@@ 4.2. Speech patterns likely to yield hallucinations
In our experiment, we compare audio from American English speakers with aphasia (5,335 audio segments) to audio spoken by the control group (i.e., those speaking “standard” American English without speech impairments, in 7,805 audio segments). First, we find that Whisper transcriptions of aphasia speakers yield significantly more hallucinations (p=0.019) than for the control group. Specifically, 1.7% aphasia audio segments yielded hallucinations, versus only 1.2% of control group audio segments (per Figure 1(b)).
@@
Next, we seek to determine if longer pauses (disfluencies) in aphasia speech could contribute to the higher rate of hallucinations. We quantify these disfluencies by using VAD (Voice Activity Detection) to measure the pauses or weak signals in audio files. First, we calculate the “non-vocal duration” of each audio file in milliseconds using PyAnnote [^7] We also run VAD analysis via PyTorch Hub using Silero (Team, 2021) and find comparable results; see the Appendix for more details. (Bredin and Laurent, 2021; Plaquet and Bredin, 2023). Then, we can calculate the share of the audio segment that is non-vocal—i.e., the percentage of audio that does not consist of any (or only very weakly detected) human speech. Per Figure 3, there are larger non-vocal shares of total durations for aphasia speakers relative to control speakers (41% versus 15%, respectively; p-value ¡ 2.2e-16). Furthermore, we can compare the non-vocal durations within each group between audio files that yielded hallucinations and audio files that did not yield hallucinations. Audio files yielding hallucinations tended to have higher non-vocal shares of durations (29% versus 26% of total durations). [^8] In Appendix Figure 5(b), we also see that (as expected), there are significantly longer non-vocal durations (in seconds) for aphasia speakers relative to control speakers (6.8 seconds versus 1.3 seconds, respectively; p-value ¡ 2.2e-16). The non-vocal durations in audio files were also significantly higher in segments yielding hallucinations relative to those that did not (5.3 seconds versus 3.5 seconds).
@@
Figure 3. Speakers with aphasia had audio files with significantly longer shares of non-vocal sounds (i.e., PyAnnote non-vocal duration in seconds, divided by total duration in seconds) relative to their control speaker counterparts. Furthermore, non-vocal shares of audio files were significantly higher for files with Whisper hallucinations as opposed to files that did not yield hallucinations. Mean non-vocal shares for aphasia speakers with hallucinations, aphasia speakers without hallucinations, control speakers with hallucinations, and control speakers without hallucinations are: 42.4%, 40.6%, 16.2%, and 15.4%, respectively.
Finally, we train a logistic regression on the audio segments with corresponding speaker demographic information to predict whether an output will be hallucinated based on speaker demographics and audio characteristics; per Table 2, we find that the most significant attributes positively corresponding to a Whisper hallucination are having longer non-vocal durations (as is the case for aphasia speakers on average), and speaking with more words. This is consistent with our hypothesis on disfluencies, and is robust to other regression specifications, including on Mahalanobis-matched subsets of audio segments (see Appendix).
Table 2. By fitting a logistic regression conditioned on speaker demographics and audio segment attributes, we find that speech with longer non-vocal periods in an audio recording, and speech with more words, tend to result in a significantly higher likelihood of a hallucinated Whisper transcription.
@@
Share of Duration Being Non-Vocal 0.951∗∗ (0.438)
Number of Words 0.056∗∗∗ (0.011)
Has Aphasia 0.368 (0.204)
Is Female --0.017 (0.168)
Age 0.044 (0.043)
@@
African American --0.132 (0.469)
Other Race --0.281 (0.518)
Years of Education --0.058 (0.033)
English is First Language --0.227 (1.035)
No Vision Loss 0.395 (1.026)
No Hearing Loss 0.500 (1.028)
Constant --6.447∗∗∗ (2.175)
@@
Log Likelihood --783.942
Akaike Inf. Crit. 1,593.883
Note: p<<<0.1; ∗∗p<<<0.05; ∗∗∗p<<<0.01
@@ 5. Implications and Future Work
Our work demonstrates that there are serious concerns regarding Whisper’s inaccuracy due to unpredictable hallucinations. It also highlights that there are significant issues with the reproducibility of transcriptions in light of non-deterministic hallucinations. Most importantly, however, there are implications for understanding the harms and biases arising from the disparate impacts on subpopulations whose speech is likely to yield hallucinations (Papakyriakopoulos et al., 2023). As shown above, there are specific harms that are produced by hallucinations disproportionately for individuals with aphasia: the perpetuation of violence, the incorrect identification or association with someone, and the false video-based authority. All three categories can have serious implications for the aphasic speaker. Hallucinated violent speech may be flagged by institutions in power, potentially leading to a premature dismissal of an applicant in a hiring process, or an individual interacting with a government service chat bot. Inaccurate identification or association, similarly, could have severe impacts on an aphasic speaker, for example in the context of legal proceedings or in interactions with insurance companies. Lastly, false authority or phishing hallucinations could result in AI-based spam detection tools misclassifying an aphasic speaker as spam, or could facilitate future cyberattacks on transcript readers.
These three types of harms can be interpreted as both allocative and representational harms (Barocas et al., 2023), i.e., harms that are caused when a system withholds from certain groups an opportunity or a resource, and harms that occur when systems reinforce the subordination of some groups along the lines of identity (including race, class, gender, etc.). This is because hallucinations can affect how individuals with aphasia can gain access to opportunities (such as jobs) and resources (such as information about government services), and because hallucinations build on existing harmful stereotypes and reinforce them.
There are additional legal and discrimination consequences to consider. Consider the use case of AI systems used for making hiring decisions. Such systems commonly involve applying speech-to-text services to transcribe video interviews, which are used by an AI tool to infer how qualified a candidate is for a job. Often companies use this score to rank candidates and reject applications. In 2023, New York City passed Local Law 144 (The New York City Council, 2021) requiring fairness audits of such hiring systems. Nationally, the Americans with Disabilities Act (ADA) (US Department of Labor, 1990) protects individuals from being unfairly evaluated in regards to their disabilities, including their speech patterns, so using biased speech-to-text systems for hiring may violate the ADA (US Equal Employment Opportunity Commission, 2008). These laws can serve as guardrails to ensure that individuals with aphasia are not disadvantaged in the hiring arena on the basis of their disability status, and serve as additional motivation to study the ways in which individuals’ disabilities might interact with the speech-to-text technology being used on their voices.
@@
Future work in this area should, therefore, examine the intersectional nature of hallucination bias and harms. We suggest that researchers may use our approach for identifying and categorizing hallucination harms, i.e., explicitly account for the type of technology (in this case: generative AI), the particular application (in this case: speech-to-text transcription), and the particular subpopulation (in this case: individuals with aphasia). This approach can avoid generalizations of AI harms and allow for more specific AI harms classifications that may lead to targeted mitigation strategies. We also suggest to decidedly include the subpopulation (in this case, people with speech impairments) into the process of designing and testing generative AI systems.
The primary outcome of this paper focuses on the existence of hallucinations, including hallucinations of different types. However, future work can and should be done using the Word Error Rate (WER) (Klakow and Peters, 2002), generally reported as a numeric percentage, which serves as the standard metric of accuracy for speech-to-text. Prior work has found that the WER metric is worse for populations with health concerns, such as individuals with dysphonia (Lopez et al., 2023). Definitionally, hallucinations will also cause the WER metric to be worse. That said, our focus on hallucinations is instructive: looking solely at the standard reported metric of WER can conceal the concrete harms of more granular text-based errors. Hallucinations can be quoted and attributed to speakers in ways affecting their employment, education, etc. more viscerally than mistranscriptions. Reading hallucinated quotes can permanently change one’s impression of the speaker in a way that simply isn’t true for a basic mistranscription (wherein a reader could easily ascertain that, e.g., “orchestra violence” refers to “orchestra violins”), even if both types of errors would lead to comparable WERs. The importance of separating WER reporting from more granular analysis of hallucinations is corroborated by contemporaneous work finding that WER “cannot differentiate between hallucinatory and non-hallucinatory models” (Frieske and Shi, 2024). Forthcoming work regards comparing WERs across different speech-to-text services, wherein we find that Whisper performs in line with expectations (Radford et al., 2022). This leads to an open question: should users accept using an on-average accurate system, in exchange for a 1% chance of yielding a potentially harmful transcription error via hallucinations?